OpenStudy (anonymous):

Find the area of the region bounded by y= x(x-1)(x-2) and the x-axis.

OpenStudy (anonymous):

I can see on my calculator that it has a region from 0 to 1 which is positive and 1 to 2 that is negative, do I add both of the areas together?

OpenStudy (shamil98):

What are the intervals?

OpenStudy (shamil98):

\[\int\limits_{a}^{b} x(x-1)(x-2) ~dx\] need the values of b and a to calculate the area.

zepdrix (zepdrix):

I hate questions like this. There are two regions, I wish they would make questions like this more clear.. I dunno what they want here..

zepdrix (zepdrix):

So you either want:\[\Large\bf\sf\int\limits\limits_{0}^{1} x(x-1)(x-2) \;dx\] or: \[\Large\bf\sf\int\limits\limits\limits_{0}^{1} x(x-1)(x-2) \;dx\quad-\quad \int\limits\limits\limits_{1}^{2} x(x-1)(x-2) \;dx\]

zepdrix (zepdrix):

I'm not sure which :( Do you have an answer key you can check?

OpenStudy (anonymous):

i do not have an answer key. but i also think that you have to add the two regions

OpenStudy (shamil98):

\[\large \int\limits_{0}^{1} x(x-1)(x-2) dx + \int\limits_{1}^{2} x(x-1)(x-2) dx\] Do you know the fundamental theorem of calculus, or are you supposed to construct a graph and use shapes to approximate the area?

OpenStudy (campbell_st):

if they are after the total region, the area below the axis will be negative and above will be positive so for mine its \[\int\limits_{0}^{1} f(x) dx + | \int\limits_{1}^{2} f(x) dx|\] hope it helps

OpenStudy (campbell_st):

so for negative area use have the integral inside absolute value symbols... that way you will be summing 2 positives..